A Complete Bibliography of Publications in Applied Mathematics and Computation: 1975–1979

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

20 December 2023 Version 1.04

Title word cross-reference

$16.00 [Ric77]. A [Mur77]. α [MH77]. β [MH77]. r [GDD77].

* [RMH77].

-distributed [GDD77]. -stable [Mur77].

0-1 [TW79].

1975 [Sco75].

4 [Ano78b].

abstract [EL77]. Academic [Sco75]. Adams [SG75]. adaptive [Tes78]. Addison [Tes77]. Addison-Wesley [Tes77]. adjoints [Gol79]. air [AMS77, Tau78]. algebraic [KM79]. Algorithm [MCRW77]. algorithms [CK79, CS76, SV75, SW79b]. Alternative [CL78]. ambient [AMS77]. analogy [ZZ76]. Analysis [MW75, AM77, AMS77, Kim76, Kol76, Tes77]. applicable [ST76]. Application [Als79, Fym79, Ser79, Tau78]. Applications [WRS77, Ano78b, HT77, Mus78]. Applied [Ano78b, Mus77, Ric77, Tes77, KK78, McG76]. approach [Lic78, LM78, Tes78]. approximate [CK79]. approximating [ST76]. Approximation [VG79, Bow78]. approximations [Ser79]. arising [EM79]. arms [GKKM77]. art [SW79a]. articular [EM79]. associated [BN75]. Asymptotic [EM79, VG79]. automated [AMS77]. autoregressive [Bre76].

Banach [LS78]. Bayes [TW79]. Beavers [Bar78]. behavior [Bre76, ST76, VG79]. Belkin [KK77]. Bellman [Sco75]. bending [ZZ76]. between [RMH77, Ste75]. bifurcation [KM79]. Blumstein [KK77]. board [Ano75b, Ano76, Ano78a]. Book [Ric77, Sco75, Tes77]. both [MH77]. boundary [Als79, BKM79, Gol75b, KM79, SV75]. boundary-value [KM79, SV75]. brain [AM77]. buckling [KSZ75].

C [Ano78b]. calculation [Str77]. care [YDIK75]. cartilage [EM79]. case [Gol75b, MW75]. Cauchy [Gol76, Gol77, HT77, HT78, Ste75]. choice [Bre76]. Chris [Tes77]. circuitry* [WRS77]. claim [VG79]. class [McG76]. clinical [MW75]. close [HKK^{+}79]. closed [EL77]. code [SW79a]. codes [SG75, Sha77]. Comments [Bar78]. communications [HKK^{+}79]. Company [Tes77]. comparing [MH77]. comparison [SV75]. compound [VG79]. Comput [Ano78b]. Computation [Tes77, Gol77]. computational [Ano78b, GKKM77, Mus77, Mus78]. computations [Bar78, DB76]. Computer [Gre78]. concepts [Ano78b, Mus77, Mus78]. conditions [Als79]. conduction [Als75]. cones [EL77]. Connection [Ste75]. consideration [MH77]. constant [MK75]. constants [BL78]. constraints [CM76]. constraints* [MCRW77]. consumption [Kim76]. control [Cas76, KY77, Sha77, Tes78]. conversion [Gol76, Gol77]. Cooke [Fym79]. Cooperative [HKKT79]. Corrections [Ano78b]. countably [Pal78]. criminal [KK77]. Critical [BN75]. cumulative [Bre76].

D [Bar78]. data [HKKN78, KY77, SW79b]. decision [RM78]. decision-node [RM78]. Degraded [HKK^{+}79]. Denman [Bar78]. dependent [MW75]. depending [KS75]. Derivatives [Vah76]. design [KY77]. Determination [MK75, MH77]. Diagonalization [Str77]. difference [Ahm79]. differential [BL78, Gla78, HKKN78, LL78b, Mur77, Ser79, Sha77]. diffusion [EM79]. Digital [RM78]. direct [ZZ76]. Discrete [KY77, MK75]. discriminant [TW79]. Distefano [Sco75]. distributed [GDD77, RMH77]. distribution [BU77, Cra79, HT77, HT78]. distributions [VG79]. diversity [BU77]. divisibility [GDD77]. dominant [Vah76]. dynamic [HKKT79]. Dynamical [SW79b]. dynamics [HKK^{+}79].

economic [HKKN78]. Ed [Ric77].

Editorial

[Ano75a, Ano75b, Ano76, Ano78a]. effects [AMS77]. Efficiencies [HT78]. efficient [CM76]. EHB [SW79b]. eigenvalue [Gla78]. ejection [YDB79]. Energy [Ric77]. Engineering [Sco75]. environmental [Fym75]. epidemiological [ST76]. equality [CM76]. equation [BN75, HKKN78, KSZ75]. equations [Ahm79, BL78, BH77, Bow78, Fym75, GKKM77, Gol76, Gol77, Gol79, KM79, LL78b, Lor78, McG76, Mur77, Ser79, Sha77]. equilibrium [Kim76]. equivalent [Gol76, Gol77]. Errata [Ano78b]. error [MH77, SG75, Sha77]. Essential [Tes77]. estimates [HT77, HT78]. Estimating [Cra79]. Estimation [KK77]. Evaluation [KM79, ZZ76]. events [MW75, RMH77]. Existence [Lor78, LL78a]. expected [ST76]. experimental [AMS77]. externalities [Kim76].

false [MH77]. Fast [KK78]. feedback [Cas76, MK75]. fields [Cas76]. filtering [RM78, Tes78]. finding [Lic78]. first [Fym75, Gol79]. first-kind [Fym75]. Fitting [HKKN78, GKKM77]. Fixed [EL77]. fluid [Gre78]. force [ZZ76]. Form [CL78]. formation [BKM79]. formula [Fym79]. formulae [VG79]. formulas [Sha76]. Fourier [Ste75]. Fred [Ric77]. Fredholm [BH77, Gol76, Gol77]. freholm [McG76]. function [Bar78, DB76]. functional [Gol75b]. Further [Ano78b, Mus78].

G [Sco75]. gains [MK75]. Galerkin [Bow78]. Gaussian [McG79]. general [Als79, Kim76]. generalization [Fym79]. generalized [Str77]. Gini [BU77]. Givens [KK78]. Glass [KK77]. Gradient [MCRW77]. Gradient-Restoration [MCRW77].

Hadamard [Tau78]. having [VG79]. health [Kim76, YDIK75]. health-consumption [Kim76]. heat [Als75]. heating [RS76]. Hermite [Mur77]. homogeneous [KK78]. hospital [YDIK75]. Hydrodynamic [ZZ76]. Hypernumbers [Ano78b, Mus77, Mus78].

II [Ano78b, Gol75b, Gol77, HKK^{+}79, MH77, Mus78]. II. [SW79a]. Ill [Gol79]. Ill-posed [Gol79]. imbedding [Als75, Als79, Gol75a, McG76, QHN78, SV75, Sco75]. index [BU77]. Industrial [Ric77]. inequalities [Lic78]. inequality [MCRW77]. infinite [Gla78, GDD77]. infinitely [Pal78]. inhomogeneous [Gol75b]. initial [KS75]. initial-value [KS75]. integer [Lic78]. integral [BL78, BH77, Fym75, Fym79, Gol76, Gol77, KSZ75, KS75, McG76]. integration [BN75]. integro [BL78]. integrodifferential [Lor78]. interconnected [AM77]. interevent [RMH77]. interpolation [Mur77, SW79b]. intersection [Wan75]. intervals [RMH77]. Introduction [Sco75]. Invariant [Gol75a, McG76, Sco75, Als75, Als79, QHN78, SV75]. Inverse [CL78]. inverses [Str77]. inversion [Fym75, Fym79]. Iteration [CL78]. Iterative [CS76, QHN78].

J [Tes77]. John [Sco75]. justice [KK77].

Kalaba [Ste75]. kernels [BH77, Bow78, KS75]. kind [Fym75, Gol79, VG79]. Kutta [SW79a, Sha76].

Laplace [LRL77]. Latin [Wan75]. left [YDB79]. lengths [BN75]. levels [AMS77]. Lg [SW79b]. Lg-spline [SW79b]. limit [KM79]. linear [Ahm79, Als79, Cas76, HKK^{+}79, Lic78, MK75]. loads [KSZ75]. Local [SG75, Sha77]. location [HT77]. loss [TW79]. lung [AMS77]. Lyapunov [LL78a]. Lyapunov-monotone [LL78a].

M [Sco75]. many [Pal78]. Markets [Pal78]. Math [Ano78b]. mathematical [MCRW77]. Mathematics [Ric77, Tes77]. matrices [Ser79, Str77]. matrix [Bar78, DB76]. mean [Ahm79]. melting [RS76]. method [BL78, Bow78, CM76, GKKM77, Gol79, KSZ75, KS75, KK78, LM78, QHN78]. methodology [AMS77]. methods [Mur77]. microsimulation [YDIK75]. Milton [Tes77]. Minimal [Cas76]. model [AMS77, Bre76, YDIK75]. models [HKKN78, ST76, KK77, Ric77]. modified [Bow78]. moment [ZZ76]. monitoring [Fym75]. monotone [LL78a, LS78]. Multicollinearity [Tin75]. multipoint [Als79]. Musès [Ano78b].

N [Bar78]. N. [Sco75]. negative [MH77]. nets [AM77]. network [RM78]. Neumann [Gre78]. neural [AM77]. neuronal [WRS77]. Newton [CM76]. node [RM78]. noise [RM78]. Nonlinear [CK79, BL78, CS76, EM79, Gla78, KM79, LM78, Sco75]. nonuniformly [RMH77]. normal [Cra79]. Numerical [Kol76, BN75, BH77, Gla78].

NY [Sco75].

observations [MW75]. observed [HKKN78]. observer [KY77]. operations [HKK^{+}79]. operators [KS75, LL78a, LS78]. Optimal [HKK^{+}79, YDB79, MK75, TW79]. optimization [CM76, KK78]. order [Ser79, SG75]. ordinary [Mur77, Sha77]. orthogonal [Kol76]. output [MK75].

P [Tes77]. PA [Ric77]. pages [Tes77]. parameter [KS75]. parameters [Cra79, HT78, KK77, LM78]. particulate [Fym75]. pattern [YDB79]. phase [RS76]. Philadelphia [Ric77]. plate [ZZ76]. point [EL77, Gol75b, SV75]. points [KM79, RMH77]. Poisson [VG79]. pole [Cas76]. pole-shifting [Cas76]. pollutants [AMS77]. pollution [Fym75, Tau78]. polynomials [Kol76]. populations [ST76]. posed [Gol79]. positive [MH77]. pp [Ric77]. prerestorative [MCRW77]. prescribed [VG79]. Press [Sco75]. Probability [Tes77]. problem [BKM79, Fym75, Gla78, RS76]. problems [Als75, Als79, CS76, Gol75b, Gol76, Gol77, KM79, MCRW77, SV75, WRS77]. procedure [TW79]. process [EM79]. Processes [Sco75, VG79]. product [GDD77, McG79, Tau78]. product-type [McG79]. programming [HKKT79, MCRW77]. properties [EM79]. Publishing [Tes77]. pulse [RS76].

Quadrature [Sha76]. quadratures [McG79]. Quasilinearization [HKKN78]. queuing [WRS77].

R [Sco75]. races [GKKM77]. radiation [Als75]. randomly [AM77]. range [Gla78]. rank [Str77]. rapidly [BH77]. Rational [Ser79, Str77]. real [Lic78]. recursive [SW79b]. regulation [YDB79]. relations [Gol75b]. Remote [Fym75, CS76, Fym79]. remote-sensing [Fym79]. Representation [RMH77, Ste75]. residual [LRL77]. resolvents [Gol77]. response [RS76]. Restoration [MCRW77]. results [Tes78]. retarded [LS78]. Review [Ric77, Sco75, Tes77]. Riccati [BN75, Gol75a]. Richardson [GKKM77]. Roberts [Ric77]. Robustness [TW79]. role [YDIK75]. root [Vah76]. Runge [SW79a, Sha76].

S [Ric77]. Sample [MH77]. sampled [KY77]. sampled-data [KY77]. Sampling [BU77]. scale [HT77]. second [Ser79]. second-order [Ser79]. sector [YDIK75]. sensing [CS76, Fym79]. sequential [MCRW77]. set [RMH77]. sets [EL77]. several [SV75]. shear [ZZ76]. shifting [Cas76]. sign [Bar78, DB76]. simple [Als75]. singularity [BN75]. size [MH77]. small [ST76]. smooth [Bow78]. smoothing [CK79]. Sobolev [LL78b, Lor78]. Society [Ric77]. Solution [Als75, BH77, Gla78, SV75]. solutions [Lic78]. solving [Gol79]. Some [Gol75b, Als75, Gol79, WRS77]. Sons [Sco75]. space [LS78]. spline [SW79b]. square [Ahm79, Wan75]. stability [Ahm79, Tes78]. stable [Mur77]. Statistical [AM77]. step [MCRW77]. stiff [Mur77]. stochastic [RM78, ST76]. stochastic-network [RM78]. stochastical [GDD77]. studies [Gre78]. support [HKK^{+}79]. surface [Wan75]. Susan [Tes77]. system [KK77, Ste75, YDIK75]. systems [Bar78, DB76, KY77, Lic78, MK75, Ser79].

tactics [HKK^{+}79]. taking [MH77]. theorems [EL77]. Theory [Tes77, BKM79, Fym79, WRS77]. therapies [MH77]. three [RS76]. three-zone [RS76]. time [HKK^{+}79]. time-varying [HKK^{+}79]. times* [LRL77]. traders [Pal78]. transformations [Gol75a, KK78]. transforms [LRL77, Ste75]. trigonometric [Ser79]. truncated [Cra79]. Tsokos [Tes77]. two [Gol75b, GDD77, MW75, MH77, RS76, SV75]. two-phase [RS76]. two-point [SV75]. type [Gre78, LL78b, LS78, Lor78, McG79, MH77].

unified [Lic78]. uniqueness [Lor78]. upon [KS75]. urine [BKM79].

V [KS75]. value [Als79, BKM79, Gol75b, KS75, KM79, SV75]. variable [SG75]. variables [GDD77]. variation [BL78, LM78]. varying [BH77, HKK^{+}79]. ventricular [YDB79]. Volterra [Bow78].

waiting [LRL77]. waves [AM77]. Wesley [Tes77]. Wiley [Sco75]. Window [HT77, HT78]. Wing [Sco75]. writing [SW79a].

xxvvii [Ric77].

zeros [LL78a, LS78]. zone [RS76].

References

Ahmadi:1979:MSS

[Ahm79] Goodarz Ahmadi. On the mean square stability of linear difference equations. Applied Mathematics and Computation, 5 (3):233–241, July 1979. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300379900158.

Alspaugh:1975:SSS

[Als75] Dale W. Alspaugh. Solution of some simple problems in the conduction and radiation of heat by invariant imbedding. Applied Mathematics and Computation, 1 (3):219–232, ???? 1975. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630037590034X.

Alspaugh:1979:AII

[Als79] Dale W. Alspaugh. Application of invariant imbedding to linear multipoint boundary value problems with general boundary conditions. Applied Mathematics and Computation, 5(3):213–232, July 1979. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300379900146.

Anninos:1977:SAB

[AM77] Photios A. Anninos and V. K. Murthy. Statistical analysis of brain waves from randomly interconnected neural nets. Applied Mathematics and Computation, 3(1):85–94, ???? 1977. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300377900133.

Azen:1977:EMA

[AMS77] S. P. Azen, J. B. Margolick, and R. P. Sherwin. An experimental model and automated methodology for the analysis of the effects of ambient levels of air pollutants on the lung. Applied Mathematics and Computation, 3(2):95–102, ???? 1977. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300377900224.

Anonymous:1975:E

[Ano75a] Anonymous. Editorial. Applied Mathematics and Computation, 1(1):??, January 1975. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300375900260.

Anonymous:1975:EB

[Ano75b] Anonymous. Editorial board. Applied Mathematics and Computation, 1(1):??, January 1975. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300375900259.

Anonymous:1976:EB

[Ano76] Anonymous. Editorial board. Applied Mathematics and Computation, 2(1):C02, ???? 1976. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300376900151.

Anonymous:1978:EB

[Ano78a] Anonymous. Editorial board. Applied Mathematics and Computation, 4(1):??, ???? 1978. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630037890022X.

Anonymous:1978:ECH

[Ano78b] Anonymous. Errata: Corrections to “Hypernumbers — II. Further concepts and computational applications”, by C. Musès, Applied Math. Comput. 4, 45–66 (1978). Applied Mathematics and Computation, 4(3):281, ???? 1978. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300378900139.

See [Mus78].

Barnett:1978:CMS

[Bar78] Stephen Barnett. Comments on “The matrix sign function and computations in systems” by E. D. Denman and A. N. Beavers. Applied Mathematics and Computation, 4(3):277–279, ???? 1978. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300378900127.

See [DB76].

Boland:1977:NSF

[BH77] W. Robert Boland and R. E. Haymond. The numerical solution of Fredholm integral equations with rapidly varying kernels. Applied Mathematics and Computation, 3(1):25–38, ???? 1977. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300377900091.

Buell:1979:BVP

[BKM79] J. Buell, R. Kalaba, and D. Marsh. On a boundary value problem in the theory of urine formation. Applied Mathematics and Computation, 5(1): 55–67, February 1979. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300379900092.

Bernfeld:1978:NVC

[BL78] S. R. Bernfeld and M. E. Lord.

A nonlinear variation of constants method for integro differential and integral equations. Applied Mathematics and Computation, 4(1):1–14, ???? 1978. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300378900231.

Boland:1975:CLN

[BN75] W. Robert Boland and Paul Nelson, Jr. Critical lengths by numerical integration of the associated Riccati equation to singularity. Applied Mathematics and Computation, 1(1):67–82, January 1975. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300375900314.

Bownds:1978:MGA

[Bow78] John M. Bownds. A modified Galerkin approximation method for Volterra equations with smooth kernels. Applied Mathematics and Computation, 4(1):67–79, ???? 1978. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300378900279.

Brewer:1976:CAM

[Bre76] James K. Brewer. A cumulative autoregressive model for choice behavior. Applied Mathematics and Computation, 2(3): 189–196, ???? 1976. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300376900011.

Bhargava:1977:SDG

[BU77] T. N. Bhargava and V. R. R. Uppuluri. Sampling distribution of Gini’s index of diversity. Applied Mathematics and Computation, 3(1):1–24, ???? 1977. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630037790008X.

Casti:1976:MCF

[Cas76] J. Casti. Minimal control fields and pole-shifting by linear feedback. Applied Mathematics and Computation, 2(1):19–28, ????

1976. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300376900175.

Chan:1979:NSA

[CK79] W. K. Chan and K. S. P. Kumar. Nonlinear smoothing: approximate algorithms. Applied Mathematics and Computation, 5 (1):1–22, February 1979. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300379900067.

Conker:1978:IIA

[CL78] Robert S. Conker and Robert J.

Lambert. Inverse iteration: an alternative form. Applied Mathematics and Computation, 4(3): 265–275, ???? 1978. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300378900115.

Childs:1976:ENM

[CM76] Bart Childs and M. J. Maron.

An efficient Newton’s method for optimization under equality constraints. Applied Mathematics and Computation, 2(3): 273–282, ???? 1976. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300376900059.

Crain:1979:EPT

[Cra79] Bradford R. Crain. Estimating the parameters of a truncated normal distribution. Applied Mathematics and Computation, 5(2):149–156, March 1979. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300379900043.

Chen:1976:IAN

[CS76] Y. M. Chen and J. Surmont. Iterative algorithms for nonlinear problems in remote sensing. Applied Mathematics and Computation, 2(3):197–228, ???? 1976. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300376900023.

Denman:1976:MSF

[DB76] Eugene D. Denman and Alex N. Beavers, Jr. The matrix sign function and computations in systems. Applied Mathematics and Computation, 2(1):63–94, ???? 1976. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300376900205.

See comments [Bar78].

Eisenfeld:1977:FPT

[EL77] J. Eisenfeld and V. Lakshmikantham. Fixed point theorems on closed sets through abstract cones. Applied Mathematics and Computation, 3(2):155–167, ???? 1977. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300377900273.

Eisenfeld:1979:APN

[EM79] Jerome Eisenfeld and Van C.

Mow. Asymptotic properties of a nonlinear diffusion process arising in articular cartilage. Applied Mathematics and Computation, 5(3):187–198, July 1979. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300379900122.

Fymat:1975:RME

[Fym75] Alain L. Fymat. Remote monitoring of environmental particulate pollution: a problem in inversion of first-kind integral equations. Applied Mathematics and Computation, 1(2):131–185, ???? 1975. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300375900090.

Fymat:1979:GCI

[Fym79] Alain L. Fymat. A generalization of Cooke’s integral inversion formula with application to remote-sensing theory. Applied Mathematics and Computation, 5 (1):23–39, February 1979. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300379900079.

Goovaerts:1977:IDP

[GDD77] M. J. Goovaerts, L. D’Hooge, and N. De Pril. On the infinite divisibility of the product of two rdistributed stochastical variables. Applied Mathematics and Computation, 3(2):127–135, ???? 1977. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630037790025X.

Garfinkle:1977:CMF

[GKKM77] J. Garfinkle, H. Kagiwada, R. Kalaba, and E. McLain. A computational method for fitting Richardson equations of arms races. Applied Mathematics and Computation, 3(2):175–185, ???? 1977. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300377900297.

Gladwell:1978:NSD

[Gla78] Ian Gladwell. On the numerical solution of a differential nonlinear eigenvalue problem on an infinite range. Applied Mathematics and Computation, 4(2): 121–137, ???? 1978. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300378900164.

Golberg:1975:IIR

[Gol75a] Michael A. Golberg. Invariant imbedding and Riccati transformations. Applied Mathematics and Computation, 1(1): 1–24, January 1975. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300375900272.

Golberg:1975:SFR

[Gol75b] Michael A. Golberg. Some functional relations for two point boundary value problems — II. The inhomogeneous case. Applied Mathematics and Computation, 1(2):93–104, ???? 1975. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300375900077.

Golberg:1976:CFI

[Gol76] Michael A. Golberg. The conversion of Fredholm integral equations to equivalent Cauchy problems. Applied Mathematics and Computation, 2(1):1–18, ???? 1976. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300376900163.

Golberg:1977:CFI

[Gol77] Michael A. Golberg. The conversion of Fredholm integral equations to equivalent Cauchy problems — II. Computation of resolvents. Applied Mathematics and Computation, 3(1):39–58, ???? 1977. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300377900108.

Golberg:1979:MAS

[Gol79] Michael A. Golberg. A method of adjoints for solving some Illposed equations of the first kind. Applied Mathematics and Computation, 5(2):123–129, March 1979. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630037990002X.

Greenspan:1978:CSN

[Gre78] Donald Greenspan. Computer studies of a von Neumann type fluid. Applied Mathematics and Computation, 4(1):15–25, ???? 1978. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300378900243.

Hess:1979:OTC

[HKK^{+}79] J. Hess, H. Kagiwada, R. Kalaba, K. Spingarn, and C. Tsokos. Optimal tactics for close support operations. II. Degraded communications with linear time-varying dynamics. Applied Mathematics and Computation, 5(3):199–212, July 1979. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300379900134.

Helms:1978:FDE

[HKKN78] Jay Helms, Harriet Kagiwada, Robert Kalaba, and John Niedercorn. Fitting differential equation models to observed economic data — I. Quasilinearization. Applied Mathematics and Computation, 4(2):139–145, ???? 1978. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300378900176.

Hess:1979:CDP

[HKKT79] James Hess, Harriet Kagiwada, Robert Kalaba, and Chris Tsokos. Cooperative dynamic programming. Applied Mathematics and Computation, 5(1): 69–74, February 1979. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300379900109.

Higgins:1977:WEL

[HT77] J. J. Higgins and D. M. Tichenor.

Window estimates of location and scale with applications to the Cauchy distribution. Applied Mathematics and Computation, 3(2):113–126, ???? 1977. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300377900248.

Higgins:1978:EWE

[HT78] J. J. Higgins and D. M. Tichenor.

Efficiencies for window estimates of the parameters of the Cauchy distribution. Applied Mathematics and Computation, 4(2): 157–165, ???? 1978. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630037890019X.

Kimbell:1976:GEA

[Kim76] Larry J. Kimbell. A general equilibrium analysis of healthconsumption externalities. Applied Mathematics and Computation, 2(1):41–62, ???? 1976. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300376900199.

Kagiwada:1977:EPB

[KK77] H. H. Kagiwada and R. E. Kalaba. Estimation of parameters in Belkin–Blumstein–Glass Models of the criminal justice system. Applied Mathematics and Computation, 3(1):59–67, ???? 1977. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630037790011X.

Kowalik:1978:FGT

[KK78] J. S. Kowalik and S. P. Kumar. Fast Givens transformations applied to the homogeneous optimization method. Applied Mathematics and Computation, 4 (3):239–252, ???? 1978. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300378900097.

Kubicek:1979:ELB

[KM79] Milan Kubíček and Miloš Marek.

Evaluation of limit and bifurcation points for algebraic equations and nonlinear boundaryvalue problems. Applied Mathematics and Computation, 5 (3):253–264, July 1979. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300379900171.

Kolb:1976:NAO

[Kol76] D. Kolb. Numerical analysis and orthogonal polynomials. Applied Mathematics and Computation, 2 (3):257–272, ???? 1976. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300376900047.

Kalaba:1975:IVM

[KS75] Robert E. Kalaba and Melvin R.

Scott. An initial-value method for integral operators V. Kernels depending upon a parameter. Applied Mathematics and Computation, 1(1):25–35, January 1975. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300375900284.

Kalaba:1975:IEM

[KSZ75] R. E. Kalaba, K. Spingarn, and E. Zagustin. On the integral equation method for buckling loads. Applied Mathematics and Computation, 1(3):253–261, ???? 1975. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300375900375.

Kumar:1977:DOD

[KY77] K. S. P. Kumar and T. H. Yam. Discrete observer in the design of sampled-data control systems. Applied Mathematics and Computation, 3(2):137–154, ???? 1977. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300377900261.

Lichtman:1978:UAF

[Lic78] Zavdi L. Lichtman. A unified approach for finding real and integer solutions to systems of linear inequalities. Applied Mathematics and Computation, 4(2): 177–186, ???? 1978. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300378900218.

Lakshmikantham:1978:EZL

[LL78a] V. Lakshmikantham and S. Leela. On the existence of zeros of Lyapunov-monotone operators. Applied Mathematics and Computation, 4(2):107–119, ???? 1978.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300378900152.

Lakshmikantham:1978:STD

[LL78b] V. Lakshmikantham and M. Lord.

Sobolev type differential equations. Applied Mathematics and Computation, 4(2):147–155, ???? 1978. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300378900188.

Lord:1978:NAM

[LM78] M. E. Lord and A. Richard Mitchell. A new approach to the method of nonlinear variation of parameters. Applied Mathematics and Computation, 4(2):95–105, ???? 1978. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300378900140.

Lord:1978:EUS

[Lor78] M. E. Lord. Existence and uniqueness of Sobolev type integrodifferential equations. Applied Mathematics and Computation, 4(3):253–263, ???? 1978. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300378900103.

Lange:1977:LTR

[LRL77] K. Lange, N. Risch, and C. Louy. Laplace transforms for residual waiting times*. Applied Mathematics and Computation, 3(3): 253–264, ???? 1977. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300377900054.

Lakshmikantham:1978:ZMO

[LS78] V. Lakshmikantham and Patrick Sutherland. On the zeros of monotone operators of retarded type in a Banach space. Applied Mathematics and Computation, 4(1):81–91, ???? 1978. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300378900280.

McGrath:1976:IIA

[McG76] Joseph F. McGrath. Invariant imbedding applied to a class of freholm integral equations. Applied Mathematics and Computation, 2(2):113–160, ???? 1976. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300376900126.

McGrath:1979:GPT

[McG79] Joseph F. McGrath. Gaussian product-type quadratures. Applied Mathematics and Computation, 5(3):265–280, July 1979.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300379900183.

Miele:1977:PSS

[MCRW77] A. Miele, A. Calabro, F. Rossi, and A. K. Wu. The prerestorative step in the sequential Gradient-Restoration Algorithm for mathematical programming problems with inequality constraints*. Applied Mathematics and Computation, 3(3):265–280, ???? 1977. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300377900066.

Murthy:1977:SSD

[MH77] Vrudhula K. Murthy and L. Julian Haywood. Sample size determination for comparing two therapies taking both the α (type I or false positive) error and the β (type II or false negative) error into consideration. Applied Mathematics and Computation, 3 (2):169–174, ???? 1977. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300377900285.

Mahesh:1975:DOC

[MK75] J. K. Mahesh and K. S. P. Kumar. Determination of the optimal constant output feedback gains for linear discrete systems. Applied Mathematics and Computation, 1(3):243–252, ???? 1975. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300375900363.

Murphy:1977:HIS

[Mur77] W. D. Murphy. Hermite interpolation and A-stable methods for stiff ordinary differential equations. Applied Mathematics and Computation, 3(2):103–112, ???? 1977. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300377900236.

Muses:1977:AHC

[Mus77] C. Musès. Applied hypernumbers: computational concepts. Applied Mathematics and Computation, 3 (3):211–226, ???? 1977. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300377900029.

Muses:1978:HIF

[Mus78] C. Musès. Hypernumbers — II. Further concepts and computational applications. Applied Mathematics and Computation, 4(1):45–66, ???? 1978. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300378900267.

See errata [Ano78b].

Murthy:1975:ACE

[MW75] V. K. Murthy and J. M.

Weiner. Analysis of clinical events: the case of two dependent observations. Applied Mathematics and Computation, 1 (1):37–46, January 1975. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300375900296.

Pallaschke:1978:MCI

[Pal78] Diethard Pallaschke. Markets with countably infinitely many traders. Applied Mathematics and Computation, 4(3):201–212, ???? 1978. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300378900073.

Quinnez:1978:IMI

[QHN78] S. Quinnez, P. Hillion, and G. Nurdin. Iterative method for invariant imbedding. Applied Mathematics and Computation, 4 (3):213–237, ???? 1978. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300378900085.

Richardson:1977:BRB

[Ric77] Harry W. Richardson. Book review: Energy: Mathematics and models: Roberts, Fred S. (Ed.), Society for Industrial and Applied Mathematics, Philadelphia, PA, 1976, xxvvii + 276 pp., $16.00. Applied Mathematics and Computation, 3(2):187, ???? 1977. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300377900303.

Roffman:1978:DFS

[RM78] Martin M. Roffman and Lawrence P. McNamee. Digital filtering of stochastic-network decision-node noise. Applied Mathematics and Computation, 4(3):189–199, ???? 1978. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300378900061.

Richardson:1977:RII

[RMH77] John Richardson, Vrudhula K. Murthy, and L. Julian Haywood. Representation of the interevent intervals between a set of nonuniformly distributed “points” or “events”*. Applied Mathematics and Computation, 3 (3):247–251, ???? 1977. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300377900042.

Rivard:1976:TPT

[RS76] J. B. Rivard and M. R. Scott. The two-phase, three-zone melting problem — response to pulse heating. Applied Mathematics and Computation, 2(2): 173–187, ???? 1976. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630037690014X.

Scott:1975:BRD

[Sco75] M. R. Scott. Book review: N.

Distefano, Nonlinear Processes in Engineering, Academic Press, NY, 1974 and R. E. Bellman and G. M. Wing, An Introduction to Invariant Imbedding, John Wiley and Sons, NY, 1975. Applied Mathematics and Computation, 1(3):279–280, ???? 1975. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300375900399.

Serbin:1979:RAT

[Ser79]  Steven M. Serbin.  Rational

approximations of trigonometric matrices with application to second-order systems of differential equations. Applied Mathematics and Computation, 5(1): 75–92, February 1979. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300379900110.

Shampine:1975:LEV

[SG75] L. F. Shampine and M. K. Gordon. Local error and variable order Adams codes. Applied Mathematics and Computation, 1 (1):47–66, January 1975. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300375900302.

Shampire:1976:QRK

[Sha76] L. F. Shampire. Quadrature and Runge–Kutta formulas. Applied Mathematics and Computation, 2 (2):161–171, ???? 1976. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300376900138.

Shampine:1977:LEC

[Sha77] Lawrence F. Shampine. Local error control in codes for ordinary differential equations. Applied Mathematics and Computation, 3(3):189–210, ???? 1977. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300377900017.

Smith:1976:AEB

[ST76] Truett L. Smith and Chris P. Tsokos. On approximating the expected behavior of stochastic epidemiological models applicable to small populations. Applied Mathematics and Computation, 2(2):95–111, ???? 1976. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300376900114.

Stenger:1975:CBC

[Ste75] Frank Stenger. Connection between a Cauchy system representation of Kalaba and Fourier transforms. Applied Mathematics and Computation, 1(1): 83–91, January 1975. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300375900326.

Strintzis:1977:DRC

[Str77] Michael G. Strintzis. Diagonalization, rank calculation and generalized inverses of rational matrices. Applied Mathematics and Computation, 3(1):69–84, ???? 1977. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300377900121.

Scott:1975:CSI

[SV75] Melvin R. Scott and Walter H.

Vandevender. A comparison of several invariant imbedding algorithms for the solution of twopoint boundary-value problems. Applied Mathematics and Computation, 1(3):187–218, ???? 1975. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300375900338.

Shampine:1979:AWR

[SW79a] Lawrence F. Shampine and Herman A. Watts. The art of writing a Runge–Kutta code. II. Applied Mathematics and Computation, 5(2):93–121, March 1979. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300379900018.

Sidhu:1979:DRA

[SW79b] Gursharan S. Sidhu and Howard L. Weinert. Dynamical recursive algorithms for Lg-spline interpolation of EHB data. Applied Mathematics and Computation, 5 (2):157–185, March 1979. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300379900055.

Tauber:1978:AHP

[Tau78] Selmo Tauber. An application of the Hadamard product to air pollution. Applied Mathematics and Computation, 4(2):167–176, ???? 1978. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300378900206.

Tesfatsion:1977:BRB

[Tes77] Leigh Tesfatsion. Book review: Probability Theory With the Essential Analysis: J. Susan Milton and Chris P. Tsokos, Applied Mathematics and Computation, Addison-Wesley Publishing Company, Inc. 1976, 335 pages. Applied Mathematics and Computation, 3(3):281, ???? 1977.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300377900078.

Tesfatsion:1978:NAF

[Tes78] Leigh Tesfatsion. A new approach to filtering and adaptive control: stability results. Applied Mathematics and Computation, 4(1):27–44, ???? 1978. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300378900255.

Tintner:1975:M

[Tin75] Gerhard Tintner. Multicollinearity. Applied Mathematics and Computation, 1(3):233–242, ???? 1975. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300375900351.

Tsokos:1979:RBO

[TW79] Chris P. Tsokos and R. L. W.

Welch. Robustness of the Bayes optimal discriminant procedure with 0-1 loss. Applied Mathematics and Computation, 5(2): 131–148, March 1979. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300379900031.

Vahrenkamp:1976:DDR

[Vah76] Richard Vahrenkamp. Derivatives of the dominant root. Applied Mathematics and Computation, 2(1):29–39, ???? 1976. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300376900187.

VanGoethem:1979:AFC

[VG79] P. Van Goethem and M. J. Goovaerts. Approximation formulae for compound Poisson processes for a kind of claim distributions having a prescribed asymptotic behavior. Applied Mathematics and Computation, 5(3):243–251, July 1979. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630037990016X.

Wang:1975:LSI

[Wan75] Alan P. Wang. Latin square and intersection of surface. Applied Mathematics and Computation, 1 (3):263–277, ???? 1975. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300375900387.

Ward:1977:AQT

[WRS77] D. S. Ward, I. Rubin, and R. J. Sclabassi. Applications of queuing theory to some problems in neuronal circuitry*. Applied Mathematics and Computation, 3 (3):227–245, ???? 1977. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300377900030.

Yamashiro:1979:ORL

[YDB79] S. M. Yamashiro, J. A. Daubenspeck, and F. M. Bennett. Optimal regulation of left ventricular ejection pattern. Applied Mathematics and Computation, 5 (1):41–54, February 1979. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300379900080.

Yett:1975:MMH

[YDIK75] Donald E. Yett, Leonard Drabek, Michael D. Intriligator, and Larry J. Kimbell. A microsimulation model of the health care system: the role of the hospital sector. Applied Mathematics and Computation, 1(2):105–130, ???? 1975. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300375900089.

Zagustin:1976:HAD

[ZZ76] Anatol I. Zagustin and Elena A.

Zagustin. Hydrodynamic analogy for direct shear force and bending moment evaluation in a plate. Applied Mathematics and Computation, 2(3):229–255, ???? 1976. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300376900035.